OpenStudy (anonymous):

Barbara draws pens randomly from a box containing 5 pens of the same shape and size. There is 1 green pen, 3 red pens, and 1 blue pen. She draws 1 red pen and then another red pen without replacing the first one. Find the probability of drawing 1 red pen followed by another red pen, and show the equation used.

OpenStudy (anonymous):

@nick1234567 @mathmath333 @jessicabraham2000

OpenStudy (anonymous):

probability=favorable cases/possible cases=3/5

OpenStudy (anonymous):

so the answer is 3/5?

OpenStudy (anonymous):

3/5= 6%

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you guys!

OpenStudy (anonymous):

That's not the answer, we need to continue...

OpenStudy (anonymous):

oh, ok. what else do we need to do?

OpenStudy (anonymous):

oh its not sorry i thought it was nvm

OpenStudy (anonymous):

Probability= Favorable Cases/Possible Cases= 2/4=...

OpenStudy (anonymous):

50%

OpenStudy (anonymous):

Good. Now.. 60/100*50/100=...

OpenStudy (anonymous):

0.3

OpenStudy (anonymous):

Great!!! 0.3=30%

OpenStudy (anonymous):

So is 30% the answer?

OpenStudy (anonymous):

0.3 cause 100/60=0.6 and 100/50 = 0.5 wich is 0.3 move dicemal point and u get 30%

OpenStudy (anonymous):

Yes! 30% is the requested probability!

OpenStudy (anonymous):

One more question? how would i put that into complete sentences?

OpenStudy (anonymous):

the percentage would be 30%.

OpenStudy (anonymous):

go from that

OpenStudy (anonymous):

Then describe the steps taken to reach that answer!!

OpenStudy (anonymous):

yes! @jessicabraham2000

OpenStudy (anonymous):

so wait how did we get 2/4

OpenStudy (anonymous):

2/4= 50% is the Probability= Favorable Cases/Possible Cases

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

Your Welcome! Feel free to give me a medal and follow me!

OpenStudy (anonymous):

The probability is 30%. I started of by using the fraction 3/5 which is 6% and then 2/4 which is 50%. Then i multiplied 3/5 and 2/4 which was .3 and then i changed it to a percentage which is 30%

OpenStudy (anonymous):

thats what i got.

OpenStudy (anonymous):

i meant off not of btw

OpenStudy (anonymous):

Looks good, but you might want to add why we used the fractions 2/4 and 3/5. Add that it is Probability= Favorable Cases/Possible Cases

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

Of Course!